OpenStudy (anonymous):

Using the given zero, find one other zero of f(x). 2 - 4i is a zero of f(x).= x4 - 4x3 + 21x2 - 4x + 20

OpenStudy (amoodarya):

as the coefficient are real numbers then another root is 2+4i

OpenStudy (amoodarya):

(x-2+4i)(x-2-4i)=x^2-4x+20 so divide f(X) to x^2-4x+20 you have degree 2 then you can do it your own try it hint :f has 4 unreal root

OpenStudy (amoodarya):

f(x)= (x2-4x+20)(x2+1) so the roots are x= 0+i , 0-i , 2-4i , 2+4i

OpenStudy (amoodarya):

is anybody there ?